OpenStudy (melstutes):

Point a is( -3, 4) and point b is (5, -2). How many units long is line segment ab. Find the ordered pair that is 1/3 of the distance of ab from endpoint a.

OpenStudy (mayankdevnani):

use DISTANCE FORMULA

OpenStudy (melstutes):

I did that and got 10

OpenStudy (mayankdevnani):

you are correct !

OpenStudy (melstutes):

I know to find distance I would divide by 3 but do not understand how to find coordinate pair 1/3 away from a

OpenStudy (mayankdevnani):

1/3 of the distance of AB=1/3 * 10=10/3

OpenStudy (melstutes):

I understand that but what is ordered pair -I am lost here

OpenStudy (mayankdevnani):

ordered pair means you have to find the other point which is 10/3 distance away from end point A. Simply,you have to calculate another point in the form of (x,y)

OpenStudy (melstutes):

I don't understand how to do that

OpenStudy (mayankdevnani):

apply distance formula for the same !

OpenStudy (mayankdevnani):